A train started from rest and moved with constant acceleration. At one time it was travelling at 33 m/s and 160m farther on it was travelling 54m /s

Calculate (a) acceleration (b) time required to travel 160m(c)distance moved from rest to time the train had a speed of 33m/s

1
Expert's answer
2022-01-06T09:07:19-0500

Explanations & Calculations


  • Since the acceleration is uniform, you can use any of the 4 motion equations appropriately.
  • Thing to take into consideration is that the initial velocity is zero.

a).

  • For this you can use the formula "\small v^2 = u^2+2as" from the starting point to the point 160m away, as you are given the values of all variable except "a" which is the acceleration you need to find.
  • Use the units correctly and get the answer.


b).

  • For this you can use "s=ut+\frac{1}{2}at^2" from the start to the point 160m away. You have now found the value of acceleration and that could be used here to calculate the time.


c).

  • For this you can use "v^2 = u^2+2as" from start to the point where the train had 33 m/s speed as you now know the values of all the variables except "s" which is the displacement that you need to calculate.



  • You now can probably give it a try.